OpenStudy (mathmath333):

whats the last digit for (2^123!)*(7^76!)

OpenStudy (anonymous):

The ones digit for powers of 2 repeats: \[2^1=2\\ 2^2=4\\ 2^3=8\\ 2^4=16\\ 2^5=32\\ 2^6=64\] and so on, which gives a pattern of \(2-4-8-6-2-4-8-6-\cdots\). The ones digit for powers of 7 also repeats: \[7^1=7\\ 7^2=49\\ 7^3=343\\ 7^4=2401\\ 7^5=16807\\ 7^6=117649\] and so on, with a pattern of \(7-9-3-1-7-9-3-1-\cdots\).

OpenStudy (anonymous):

The only foreseeable problem is checking the value of 123!... Most scientific calculators aren't equipped for numbers of that magnitude...

ganeshie8 (ganeshie8):

also u may work mod 10 directly : last digit = (2^123!)*(7^76!) mod 10

ganeshie8 (ganeshie8):

this may help in reducing : 7^2 = 49 = -1 mod 10

OpenStudy (anonymous):

2^(4x) MOD 10 = 6 7^(4y) MOD 10 = 1 So, 2^(4x) * 2^(4y) MOD 10 = 6

OpenStudy (anonymous):

Since 123! and 76! can be expressed as 4k, last digit of 2^123! * 7^76! is 6

OpenStudy (mathmath333):

ok thns